Assume that when adults with smartphones are randomly selected, 44% use them in meetings or classes. If 8 adult smartphone users are randomly selected, find the probability that at least 5 of them use their smartphones in meetings or classes. 

The probability is _______ 
Answer #1

Binomial distribution: P(X) = nCx px qn-x

n = 8

p = 0.44

q = 0.56

P(at least 5 of them use smartphones in meetings or classes) = P(X \geq 5)

= P(5) + P(6) + P(7) + P(8)

= 8C5x0.445x0.563 + 8C6x0.446x0.562 + 8x0.447x0.56 + 0.448

= 0.1622 + 0.0637 + 0.0143 + 0.0014

= 0.2416